CONFIRMATION OF VIBRATION COUPLING IN THE SYMMETRIC CH STRETCH AS REVEALED BY COHERENCE-DETECTED FTMW-IR SPECTROSCOPY OF CH_{3}OD.

\underline{SYLVESTRE TWAGIRAYEZU}, TROCIA N. CLASP, DAVID S. PERRY, Department of Chemistry, The University of Akron, Akron OH 44325; JUSTIN L. NEILL, MATT T. MUCKLE, and BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904.

Rotational state-selection with E-species transitions (3_{0} ← 3_{−1},2_{0} ← 3_{1},1_{0} ← 1_{1}) is employed to record the infrared spectra in the range 2750-2900 cm^{−1}. The observed spectra of CH_{3}OD contain only one vibrational band origin whereas 12 interacting vibrational bands of CH_{3}OH are observed in the same interval. There are no interacting vibrational bands in the CH_{3}OD spectra because the first tier states are shifted out of resonance. This observation confirms that the dominant coupling pathway in the ν_{3} CH stretch of CH_{3}OH is the third order coupling of the CH stretch to a combination of the COH bend and an HCH bend, which subsequently couples to high order combination bands involving torsional excitation.